OpenStudy (anonymous):

Rewrite y = x2 + 14x + 29 in general form.

OpenStudy (anonymous):

it already is in general form

OpenStudy (anonymous):

(x+7)^2 - y -20 = 0

OpenStudy (anonymous):

^bro wtf you talkin' about. y= x2+14x+29 is in general form

OpenStudy (anonymous):

y = (x+7)^2 - 20 is the vertex form

OpenStudy (anonymous):

It's not a linear equation, so I don't see any general form here

OpenStudy (anonymous):

It tells me to rewrite the Standard Form of a Quadratic Equation in General Form

OpenStudy (anonymous):

darlin' that is in general form

OpenStudy (anonymous):

y-29=x^2+14x y-29+14=x^2+14x+14 y-29+14=(x+7)^2 y-15=(x+7)^2 y=(x+7)^2 y=1(x+7)^2+15 what is wrong